Temperature, x :99 87 81 91 95 86 94 96

Number of absences, y :14 6 3 6 10 5 7 11

1. Find the equation of the regression line. (Round to 2 decimals)

2. Find the correlation coefficient r. (Round to 3 decimals)

3. State the strength and direction of the correlation that exists.
Answer #1

1)

х Y X*Y X2 y2 99 14 1386 9801 196 87 6 522 7569 36 81 3 243 6561 9 91 6 546 8281 36 95 10 950 9025 100 86 5 430 7396 25 94 7

2)

SSxy = 5791 (72 (729 x 62) = 141.25 1 SSxx = 66685 (729)2 = 254.875 8 SSyy = 572 (62)2 = (62)2 = 91.5 8 Therefore, based on t

3) 0.925 is very close to 1 .it means there is a strong and positive correlation between variables

please like ??